OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, -4).

OpenStudy (anonymous):

these are the answer choices; y =-1/4 x2 y2 = -4x y2 = -16x y = -1/16 x2

OpenStudy (anonymous):

how do you know?